OpenStudy (anonymous):

Each year, Terrell buys a plane ticket to visit his grandmother for a week. Last year, he paid . This year, he paid . What was the percent increase in the cost of his plane ticket? Show your work

OpenStudy (anonymous):

143-130=13 13 = 0.01*x*130 13 = 1.3*x 10 = x (10% increase)